OpenStudy (anonymous):

write the slope -intercept equation for the line with slope -3/7 and y-intercept (0,-5)

OpenStudy (anonymous):

You need to put this into point-slope. y-y1=m(x-x1). y+5=-3/7(x-0) y+5=-3/7x+0 Now we subtract 5 on both sides and we get y=-3/7x-5 and it's in slope-intercept form.